I think this problem mostly has to do with the discriminant,but I tried the ways showed on the website and I'm still confused.
If 4x^2+kx+6=0 has one real root, find k.

Answer by stanbon(75887)

If 4x^2+kx+6=0 has one real root, find k.
----------------
You are right about the discriminant.
The discriminant must be zero if there is to be only one real root.
-------------------
k^2-4*4*6=0
k^2-96=0
k = sqrt96 or k=-sqrt96
k = 4sqrt6 or k= -4sqrt6
===========================
Cheers,
Stan H.